# Stabilizing the Splits through Minimax Decision Trees

### v2

@@ 2.3 Empirical risk bound for cyclic MinimaxSplit
So far, we have discussed the cases where the joint distribution $(\mathbf{X},Y)$ is marginally atomless, which does not include the class of empirical measures. We now focus on the regression setting and derive finite-sample performance guarantees. Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and $Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon:=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]+\varepsilon,$ where the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$, $\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp(-{u^{2}}/({2\sigma^{2}})),\leavevmode\nobreak\ u\geqslant 0.$ Also, we consider the empirical law of $(\mathbf{X},Y)$ of $N$ samples from $\mathbb{P}_{*}$.
@@
Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
@@
Our next key result is the following oracle inequality for decision trees under model mis-specification (i.e., when $g_{*}$ does not belong to $\mathcal{G}$). Let $g_{k}$ be the output of the cyclic MinimaxSplit construction (that is, $M_{k}=g_{k}(\mathbf{X})$). Note that $g_{k}$ is random under the law $\mathbb{P}_{*}$. We use $\left\lVert\cdot\right\rVert$ to denote the $L^{2}$ norm in $\mathbb{P}_{*}$, e.g. $\left\lVert g_{k}-g_{*}\right\rVert^{2}=\mathbb{E}[(g_{k}(\mathbf{X}_{*})-g_{*}(\mathbf{X}_{*}))^{2}]$, where the expectation is taken in $\mathbb{P}_{*}$.
@@
For a fixed sample size $N$, if we assume that $g_{*}\in\mathcal{G}$ and $k=3d\log_{2}N/(3d+2)$ is a multiple of $d$, Theorem 7 has the further consequence that (under the same assumptions)
$\mathbb{E}\big{[}\left\lVert g_{k}-g_{*}\right\rVert^{2}\big{]}\leqslant CN^{-\frac{2}{3d+2}}\big{(}\left\lVert g_{*}\right\rVert_{\mathrm{TV}}+(\log N)^{2}\log(Nd)\big{)}.$ (24)
@@ 2.4 Why cyclic MinimaxSplit outperforms MinimaxSplit
This ASBP phenomenon has an important consequence of the inadmissibility of classic random forests in the context of MinimaxSplit. We follow the setting of Section 7 of Klusowski and Tian (2024), which is a slight adaptation of the Breiman’s classic random forest (Breiman, 2001). A random forest consists of a collection of independent subsampled trees, where at each node, the split dimensions are optimized among a random pre-selected set $\mathcal{S}\subseteq[d]$. The set $\mathcal{S}$ is independently randomly generated with a fixed size $m_{\mathrm{try}}\in[d]$ for each node. The notation $m_{\mathrm{try}}$ will be kept throughout this paper. Consider a random forest under the MinimaxSplit regime, with $m_{\mathrm{try}}\geqslant 2$ and $g(\mathbf{x})=x_{1}+\dots+x_{d-1}+|x_{d}|$ on $1,1]^{d}$. Then each $\mathcal{S}$ must include a dimension in [d1][d-1], along which the split always reduces more risk than the $d$-th dimension. This implies that the $d$-th dimension is never reached and the risk is strictly bounded away from zero for any $k$, using an argument similar to (25). Therefore, the consistency of the random forest under the MinimaxSplit regime is not always guaranteed in the large-sample regime (even if the signal is noiseless), unless we pick $m_{\mathrm{try}}=1$. In the next section, we analyze the case $m_{\mathrm{try}}=1$, which we call a random-dimension random forest, and establish the corresponding consistency for the MinimaxSplit random forest. The empirical performance will be presented in Section 5.3.
@@ 3.1 MinimaxSplit random forests
The ensemble method we consider here is the original Breiman’s random forest (Breiman et al., 1984), which aggregates multiple decision trees to improve the precision and robustness of the prediction. Each tree is grown to a fixed depth $k$ according to some splitting rule, and the split dimension of each node is optimized among independent selections of random subsets of $[d]$ with a fixed cardinality $m_{\mathrm{try}}$. Each tree in the ensemble is trained on a bootstrap sample of the original dataset and predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model (see Algorithm 1 in Appendix E). We call the ensemble the MinimaxSplit (resp. VarianceSplit) random forest, if the base learners follow the MinimaxSplit (resp. VarianceSplit) algorithm. [^6] Note that the cyclic MinimaxSplit case does not apply (unless $m_{\mathrm{try}}=1$) due to the lack of an optimization procedure over the dimensions.
@@ 3.2 Empirical risk bound for MinimaxSplit random-dimension random forests
Our discussions in Section 2.4 indicate that the random forest approach for the MinimaxSplit algorithms is better performed with $m_{\mathrm{try}}=1$ to guarantee consistency. In other words, we randomize the dimension at every node. As a consequence, applying the MinimaxSplit and the cyclic MinimaxSplit splitting rules yield the same ensemble in distribution. We call the resulting ensemble the MinimaxSplit random-dimension random forest.
Recall the regression setting from Section 2.3 and consider a random forest consisting of trees in which split dimensions at each node are i.i.d. sampled from $[d]$, independently of the data, similar to ExtraTrees (Geurts et al., 2006) but the randomness lies in choosing the dimension instead of the split location. Fix a depth $k$ and the number of samples $N$. Denote by $n$ the number of trees in the forest. Let $\Sigma$ denote the law of the splitting dimensions and $\mathbb{E}_{n}$ denote the expectation with respect to the empirical measure on $n$ samples from $\Sigma$. Let $g_{k}$ be the estimator constructed at deapth $k$ from the MinimaxSplit random-dimension random forest. Our goal is then to bound the $L^{2}$ error $\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]},$where the outer expectation is taken in terms of randomness both from the empirical samples from $(\mathbf{X}_{*},Y_{*})$ and from the empirical measure $\mathbb{E}_{n}$. We will provide empirical evidence of the effectiveness of such a randomised construction in Section 5.3.
@@
Assume the same regression setting of Theorem 7, and the above random forest definition. We have
$\begin{split}\mathbb{E}\big{[}\left\lVert\mathbb{E}_{n}[g_{k}]-g^{*}\right\rVert^{2}\big{]}\leqslant&\frac{C2^{k}(\log N)^{2}\log(Nd)+2^{k}(\log N)dn^{-1/3}}{N}\\ &+2\inf_{g\in\mathcal{G}}\bigg{(}2\left\lVert g-g^{*}\right\rVert^{2}+\frac{8}{3}\Big{(}\left\lVert g-g^{*}\right\rVert^{2}+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big{)}\Big{(}e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big{)}\bigg{)},\end{split}$ (26)
where $C>0$ is a constant depending only on $\left\lVert g^{*}\right\rVert_{\infty}$ and $\sigma^{2}$. In particular, if $N/(2^{k}(\log N)^{2}\log(Nd))\to\infty$, $k/d\to\infty$, and $n/d^{3}\to\infty$, then the MinimaxSplit random-dimension random forest is consistent.
@@ 4 Minimax decision trees for classification
Given a split rule, we obtain a nested sequence of axis-aligned partitions $\{\pi_{k}\}_{k\geqslant 0}$ by applying a single binary split per level. For any $\omega$, we let $A_{k}(\omega)\in\pi_{k}$ be the unique cell containing $\mathbf{X}(\omega)$. The node statistic we track is the log-odds of the average class-probability in that cell,
@@
Let $g_{k}(\mathbf{x}):=\log(\eta_{A_{k}}/(1-\eta_{A_{k}}))$ for $\mathbf{x}\in A_{k},\leavevmode\nobreak\ A_{k}\in\pi_{k}$. The predictor associated with depth $k$ is the plug-in classifier $2{\mathbbm{1}}_{\{g_{k}(\mathbf{x})\geqslant 0\}}-1$ (or equivalently, $2{\mathbbm{1}}_{\{\eta_{A_{k}}\geqslant{1}/{2}\}}-1$ for the unique $A_{k}\in\pi_{k}$ that contains $\mathbf{x}$) and the canonical convex surrogate is the conditional cross-entropy on the partition.
With this preparation, the results in this section follow the same logical arc as in the regression setting. Theorem 10 treats the marginally atomless case and shows that, under the cyclic MinimaxSplit schedule, the log-odds process $\{M_{k}\}$ contracts in mean square at an exponential rate that depends only on the dimension and the coordinate range. Theorem 11 covers the complementary purely atomic setting. Here, we need a mild anti-concentration assumption on coordinate masses; the theorem then delivers an explicit upper bound for $\mathbb{E}[(Y-M_{k})^{2}]$ that reflects the discrete resolution $N$ of the atoms. Finally, Theorem 12 translates these bounds into an oracle inequality for the excess classification risk of $g_{k}$ under the 0-1 loss via standard calibration between cross-entropy and misclassification.
@@
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big{)}.\end{split}$ (29)
@@
Suppose that $\mathbf{X}$ is purely atomic and for some $N>0$,
@@
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant 2^{-\lfloor k/d\rfloor/2}\frac{2^{k}\log N}{N}+\inf_{g\in\mathcal{G}}\Big{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2})\Big{)}.\end{split}$ (31)
The proofs of Theorems 10 and 11 follow essentially the same principles as in the regression setting, except that now $\{M_{k}\}_{k\geqslant 0}$ no longer forms a martingale. The key is to apply the monotonicity of the map $z\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}<z)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<z))$, a fact that mirrors Lemma 2.1 and will be justified in Lemma B.3 below.
In the next result, we state the corresponding oracle inequality. Following Klusowski and Tian (2024), for a map $g$ we consider the misclassification risk $\mathrm{Err}(g):=\mathbb{P}(Y\neq 2{\mathbbm{1}}_{\{g(\mathbf{X})\geqslant 0\}}-1).$ Further, we set $g_{*}(\mathbf{x}):=\log(\mathbb{P}(Y=1\mid\mathbf{X}=\mathbf{x})/\mathbb{P}(Y=-1\mid\mathbf{X}=\mathbf{x}))$.
@@
for some large constant $L>0$, using the rough inequality $x\leqslant\min\{x^{2}+1,x^{4}+1\}$. Compared to the oracle inequality of Klusowski and Tian (2024), we see that we have achieved exponential convergence rates and removed the factor $\left\lVert g\right\rVert_{\infty}$. The proof of Theorem 12 follows in exactly the same way as Theorem 4.3 of Klusowski and Tian (2024) in the C4.5 case, as the arguments therein do not depend on how the tree is constructed, as long as the splitting rule is axis-aligned.

### v3

@@ 2.3 Empirical risk bound for cyclic MinimaxSplit
So far, we have discussed the cases where the joint distribution $(\mathbf{X},Y)$ is marginally atomless, which does not include the class of empirical measures. We now focus on the regression setting and derive finite-sample performance guarantees. Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and $Y_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon,$ where $g_{*}(\mathbf{X}_{*}):=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]$ and the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$, $\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp(-{u^{2}}/({2\sigma^{2}})),\penalty 10000\ u\geqslant 0.$ Also, we consider the empirical law of $(\mathbf{X},Y)$ based on $N$ samples from $\mathbb{P}_{*}$.
@@
When the marginal law of $X$ has atoms, the continuity argument used in the proof of the exact worst-child halving property is no longer available. Nevertheless, if every coordinate atom has mass at most $1/N$, then the failure of exact balancing can be quantified by an additive perturbation of order $N^{-1}$. The next theorem shows that the population bound from the atomless case survives, up to a discrete-resolution correction term.
@@
Suppose that $\mathbf{X}$ is purely atomic and has a finite support, and for some $N>0$,
@@
Comparing (23) and (13), we see that there is an additional compensation term of $(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}(\Delta Y)^{2}}{3N}$ for the atomic case. This arises because the functions $\varphi_{L}$ and $\varphi_{R}$ are no longer continuous in Lemma 2.1, meaning that the maximum risk decay may not always achieve the $1/2$ rate; see Remark 1.
The previous results control the approximation error of the population partition generated by cyclic MinimaxSplit. To obtain a finite-sample oracle inequality, one combines that deterministic approximation bound with the same truncation and empirical-process steps used in the CART analysis of Klusowski and Tian (2024). The effect of the MinimaxSplit rule enters only through the improved approximation term.
@@
For a fixed sample size $N$, if we assume that $g_{*}\in\mathcal{G}$ and $k=\lceil(3d\log_{2}N)/(3d+2)\rceil$ is a multiple of $d$, Theorem 8 has the further consequence that (under the same assumptions)
$\mathbb{E}\big[\left\lVert g_{k}-g_{*}\right\rVert^{2}\big]\leqslant CN^{-\frac{2}{3d+2}}\big(\left\lVert g_{*}\right\rVert^{2}_{\mathrm{TV}}+(\log N)^{2}\log(Nd)\big).$ (24)
@@
As explained in Remark 4, Theorem 8 does not hold without extra assumptions if the cyclic MinimaxSplit algorithm is replaced by the MinimaxSplit algorithm. On another note, in high-dimensional settings, the MSE decay rate $2^{-2\lfloor k/d\rfloor/3}$ deteriorates under the cyclic schedule. Nevertheless, we provide numerical support in Appendix E.3 that the MinimaxSplit algorithm in high dimensions still often outperforms other baseline models such as scikit-learn and VarianceSplit.
@@ 2.4 Why cyclic MinimaxSplit outperforms MinimaxSplit
This ASBP phenomenon has an important consequence for the inadmissibility of classic random forests in the context of MinimaxSplit. We follow the setting of Section 7 of Klusowski and Tian (2024), which is a slight adaptation of Breiman’s classic random forest (Breiman, 2001). A random forest consists of a collection of independent subsampled trees, where at each node, the split dimensions are optimized among a random pre-selected set $\mathcal{S}\subseteq[d]$. The set $\mathcal{S}$ is independently randomly generated with a fixed size $m_{\mathrm{try}}\in[d]$ for each node. The notation $m_{\mathrm{try}}$ will be kept throughout this paper. Consider a random forest under the MinimaxSplit regime, with $m_{\mathrm{try}}\geqslant 2$ and $g(\mathbf{x})=x_{1}+\dots+x_{d-1}+|x_{d}|$ on $1,1]^{d}$. Then each $\mathcal{S}$ must include a dimension in [d1][d-1], along which the split always reduces more risk than the $d$-th dimension. This implies that the $d$-th dimension is never reached and the risk is strictly bounded away from zero for any $k$, using an argument similar to (25). Therefore, the consistency of the random forest under the MinimaxSplit regime is not always guaranteed in the large-sample regime (even if the signal is noiseless), unless we pick $m_{\mathrm{try}}=1$. In the next section, we analyze the case $m_{\mathrm{try}}=1$, which we call a random-dimension random forest, and establish the corresponding consistency for the MinimaxSplit random forest. The empirical performance will be presented in Section 5.3.
@@ 3.1 MinimaxSplit random forests
The ensemble method we consider here is Breiman’s classic random forest (Breiman, 2001), which aggregates multiple decision trees to improve the precision and robustness of the prediction. Each tree is grown to a fixed depth $k$ according to some splitting rule, and the split dimension of each node is optimized among independent selections of random subsets of $[d]$ with a fixed cardinality $m_{\mathrm{try}}$. Each tree in the ensemble is trained on a bootstrap sample, yet our theoretical results do not consider bootstrap for simplicity. Predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model (see Algorithm 1 in Appendix F). We call the ensemble the MinimaxSplit (resp. VarianceSplit) random forest, if the base learners follow the MinimaxSplit (resp. VarianceSplit) algorithm. [^6] Note that the cyclic MinimaxSplit case does not apply (unless $m_{\mathrm{try}}=1$) due to the lack of an optimization procedure over the dimensions.
@@ 3.2 Empirical risk bound for MinimaxSplit random-dimension random forests
Our discussion in Section 2.4 indicates that the random forest approach for the MinimaxSplit algorithms is better implemented with $m_{\mathrm{try}}=1$ to guarantee consistency. In other words, we randomize the dimension at every node. As a consequence, applying the MinimaxSplit and the cyclic MinimaxSplit rules yield the same ensemble in distribution. We call the resulting ensemble the MinimaxSplit random-dimension random forest.
Recall the regression setting from Section 2.3 and consider a random forest consisting of trees in which split dimensions at each node are i.i.d. sampled from $[d]$, independently of the data, similar to ExtraTrees (Geurts et al., 2006) but the randomness lies in choosing the dimension instead of the split location. Fix a depth $k$ and the number of samples $N$. Denote by $n$ the number of trees in the forest. Let $\Sigma$ denote the law of the splitting dimensions and $\mathbb{E}_{n}$ denote the expectation with respect to the empirical measure on $n$ samples from $\Sigma$. Let $g_{k}$ be the estimator constructed at depth $k$ from the MinimaxSplit random-dimension random forest. Our goal is then to bound the $L^{2}$ error $\mathbb{E}\big[\left\lVert\mathbb{E}_{n}[g_{k}]-g_{*}\right\rVert^{2}\big],$ where the outer expectation is taken in terms of randomness both from the empirical samples from $(\mathbf{X}_{*},Y_{*})$ and from the empirical measure $\mathbb{E}_{n}$. We will provide empirical evidence of the effectiveness of such a randomized construction in Section 5.3.
@@
Assume the same regression setting as in Theorem 8, and the above random forest definition. We have
$\begin{split}\mathbb{E}\big[\left\lVert\mathbb{E}_{n}[g_{k}]-g_{*}\right\rVert^{2}\big]\leqslant&\frac{C2^{k}(\log N)^{2}\log(Nd)+2^{k}(\log N)dn^{-1/3}}{N}\\ &+2\inf_{g\in\mathcal{G}}\bigg(2\left\lVert g-g_{*}\right\rVert^{2}+\frac{8}{3}\Big(\left\lVert g-g_{*}\right\rVert^{2}+\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\Big)\Big(e^{-\frac{k}{6d}}+\frac{d}{n^{1/3}}\Big)\bigg),\end{split}$ (26)
where $C>0$ is a constant depending only on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma^{2}$. In particular, if $N/(2^{k}(\log N)^{2}\log(Nd))\to\infty$, $k/d\to\infty$, and $n/d^{3}\to\infty$, then the MinimaxSplit random-dimension random forest is consistent.
@@
The restriction $m_{\mathrm{try}}=1$ is theoretically motivated: for $m_{\mathrm{try}}\geqslant 2$, Example 9 shows that the obstruction to extending Theorem 3 from cyclic MinimaxSplit to the unrestricted MinimaxSplit rule is not the local minimax contraction itself, but rather a coordinate-starvation phenomenon: a signal-bearing coordinate may never be selected along some root-to-leaf paths. The next condition rules this out at the level of the selected split coordinate.
Let $\Xi$ denote all randomization used by the tree construction (candidate-set sampling and tie-breaking), and let $\{\pi_{k}^{\Xi}\}_{k\geqslant 0}$ be the associated random partition sequence. For a splittable cell $A\in\pi_{k}^{\Xi}$, write $J^{\Xi}(A)\in[d]$ for the coordinate along which $A$ is split.
Definition 3.1 (Positive visibility on $S$).
Fix $S\subseteq[d]$ and $\rho\in(0,1]$. We say that the randomized MinimaxSplit tree has positive visibility $\rho$ on $S$ if for every $k\geqslant 0$, every splittable cell $A\in\pi_{k}^{\Xi}$, and every $j\in S$,
$\mathbb{P}\!\left(J^{\Xi}(A)=j\,\middle|\,\pi_{k}^{\Xi}\right)\geqslant\rho\qquad\text{a.s.}$
In the following, for $S\subseteq[d]$, we consider the additive function class
$\mathcal{G}_{S}:=\Big\{g(\mathbf{x}):=\sum_{j\in S}g_{j}(x_{j})\Big\},$
where $\mathbf{x}=(x_{1},\dots,x_{d})$.
Theorem 11.
Suppose that $\mathbf{X}$ is marginally atomless. Let $\{M_{k}^{\Xi}\}_{k\geqslant 0}$ be the predictor sequence associated with a randomized MinimaxSplit tree, and assume that the tree has positive visibility $\rho\in(0,1]$ on $S\subseteq[d]$ in the sense of Definition 3.1. Then uniformly for any $\delta>0$ and $k\geqslant 0$,
$\begin{split}\mathbb{E}[(Y-M_{k}^{\Xi})^{2}]\leqslant\inf_{g\in\mathcal{G}_{S}}\Bigg(\left((1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3}e^{-\rho k/3}\right)\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\\ (1+\delta^{-1})\left(\frac{1}{3}+\left(\frac{1+\delta^{-1}}{4}\right)^{2/3}\right)e^{-\rho k/3}\|g\|_{\mathrm{TV}}^{2}\Bigg),\end{split}$ (27)
where the expectation is taken over both $(\mathbf{X},Y)$ and the tree randomization $\Xi$.
In particular, if $\rho$ is bounded below by a positive multiple of $1/s$, where $s:=|S|$, then the approximation error decays at rate $\exp(-ck/s)$ for a universal constant $c>0$. This reveals that the stochastic visibility scale $1/\rho$ for randomized trees with $m_{\mathrm{try}}\geqslant 2$ is essential to ensure risk bounds. This kind of assumption shares the spirit of mild design assumptions used by (Ročková and Saha, 2019). The same observation also shows that randomized MinimaxSplit adapts to exact sparsity under the positive visibility condition. Indeed, as the approximation error decays at rate $\exp(-ck/s)$, the effective dimension is $s$ rather than the ambient dimension $d$.
@@ 4 Minimax decision trees for classification
Given a split rule, we obtain a nested sequence of axis-aligned partitions $\{\pi_{k}\}_{k\geqslant 0}$ by applying a single binary split per node per level. For any $\omega$, we let $A_{k}(\omega)\in\pi_{k}$ be the unique cell containing $\mathbf{X}(\omega)$. The node statistic we track is the log-odds of the average class-probability in that cell,
@@
If a cell is pure, we allow $M_{k}=\pm\infty$ in the natural way; then $\log(1+e^{-YM_{k}})=0$ on that cell. Let $g_{k}(\mathbf{x}):=\log(\eta_{A_{k}}/(1-\eta_{A_{k}}))$ for $\mathbf{x}\in A_{k},\penalty 10000\ A_{k}\in\pi_{k}$. The predictor associated with depth $k$ is the plug-in classifier $2{\mathbbm{1}}_{\{g_{k}(\mathbf{x})\geqslant 0\}}-1$ (or equivalently, $2{\mathbbm{1}}_{\{\eta_{A_{k}}\geqslant{1}/{2}\}}-1$ for the unique $A_{k}\in\pi_{k}$ that contains $\mathbf{x}$) and the canonical convex surrogate is the conditional cross-entropy on the partition.
With this preparation, the results in this section follow the same logical arc as in the regression setting. Theorem 12 treats the marginally atomless case and shows that, under the cyclic MinimaxSplit schedule, the cross entropy contracts exponentially in $\lfloor k/d\rfloor$. Theorem 13 covers the complementary purely atomic setting. Here, we need a mild anti-concentration assumption on coordinate masses; the theorem then delivers an explicit upper bound for $\mathbb{E}[\log(1+e^{-YM_{k}})]$ that reflects the discrete resolution $N$ of the atoms. Finally, Theorem 14 translates these bounds into an oracle inequality for the excess classification risk of $g_{k}$ under the 0-1 loss via standard calibration between cross-entropy and misclassification.
@@
$\begin{split}\mathbb{E}[\log(1+e^{-YM_{k}})]&\leqslant\inf_{g\in\mathcal{G}}\Big(\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})})]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big).\end{split}$ (30)
@@
Suppose that $\mathbf{X}$ is purely atomic and has a finite support, and for some $N>0$,
@@
$\begin{split}\mathbb{E}[\log(1+e^{-YM_{k}})]&\leqslant 2^{-\lfloor k/d\rfloor/2}\frac{2^{k}\log(eN)}{N}+\inf_{g\in\mathcal{G}}\Big(\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})})]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}(\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}+\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2})\Big).\end{split}$ (32)
The proofs of Theorems 12 and 13 follow essentially the same principles as in the regression setting, except that now $\{M_{k}\}_{k\geqslant 0}$ no longer forms a martingale. The key is to apply the monotonicity of the map $z\mapsto\mathbb{P}(\mathbf{X}\in A,\,X_{j}<z)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A,\,X_{j}<z))$, a fact that mirrors Lemma 2.1 and will be justified in Lemma B.5 below.
In the next result, we state the corresponding oracle inequality. Following Klusowski and Tian (2024), for a map $g$ we consider the misclassification risk $\mathrm{Err}(g):=\mathbb{P}(Y\neq 2{\mathbbm{1}}_{\{g(\mathbf{X})\geqslant 0\}}-1).$ Further, we set $g_{*}(\mathbf{x}):=\log(\mathbb{P}(Y=1\mid\mathbf{X}=\mathbf{x})/\mathbb{P}(Y=-1\mid\mathbf{X}=\mathbf{x}))$. Recall that $N$ is the number of samples, $d$ is the underlying dimension, and $g_{k}$ is the output of the cyclic MinimaxSplit construction with depth $k$.
@@
for some large constant $L>0$, using the rough inequality $x\leqslant\min\{x^{2}+1,x^{4}+1\}$. Compared to the oracle inequality of Klusowski and Tian (2024), we see that we have achieved exponential convergence rates and removed the factor $\left\lVert g\right\rVert_{\infty}$.
We provide additional classification experiments using MinimaxSplit in Appendix D.